### v1

@@ B.4 Fine-tune datasets
GeoLifeCLEF24: GeoLifeCLEF is benchmark dataset with a training set of close to 5 million plant occurrences in Europe (single-label, presence-only data) as well as a validation set of about 5000 plots and a test set with 20000 plots, with all the present species (multi-label, presence-absence data) [36]. A visual depiction of how the data look like can be found on Figure 10.
@@
CHELSA v2.1: CHELSA (Climatologies at High resolution for the Earth’s Land Surface Areas) delivers global grids of mean, minimum and maximum 2 m air temperature (tas, tasmin, tasmax) and precipitation rate (pr) at 30-arc-second resolution ($\sim 1$ km) from 1979 - present. Temperature fields are produced through lapse-rate–based statistical downscaling of ERA-5/ERA-Interim reanalyses, while precipitation is downscaled with an orographic algorithm that incorporates wind fields, boundary-layer height and subsequent bias-correction with Global Precipitation Climatology Centre (GPCC) gauge data, yielding markedly improved representation of mountain rain-shadows and thermal gradients compared with coarser products [37]. These high-fidelity temperature and precipitation layers form the foundation for CHELSA’s 19 bioclimatic derivatives and for time-series forcings such as CHELSA-W5E5, and are now a de-facto standard in species-distribution modelling, trait–environment analyses, hydrological and vegetation-dynamics models, and climate-change impact assessments. Because ecological responses to climate are often threshold-driven and spatially heterogeneous, the $\sim 1$ km detail of CHELSA allows modellers to resolve local refugia, elevational turnover and fine-scale moisture stress that coarser climatologies fail to capture, thereby increasing predictive accuracy and reducing uncertainty across taxa and regions [55].
@@ C.1 Model card
Table 6: Model card.
Model patch_size num_heads embed_dim depth swin_backbone_size Model Size
Small 4 12 384 6 medium 440M
Medium 2 16 512 10 large 980M
Table 7: Swin‐backbone configurations for Large and Medium model sizes
@@
drop_rate 0.0 0.0
attn_drop_rate 0.0 0.0
@@
Furthermore, to enhance performance and computational efficiency, particularly given the high dimensionality of the input and output data, several techniques are used. Group-Query Attention [56] is used within the Perceiver IO attention layers to reduce the memory footprint associated with key-value projections during cross-attention. Additionally, standard regularization techniques such as Layer Normalization and Dropout are applied throughout the Perceiver modules. The Swin Transformer backbone makes use of stochastic depth [57] to improve generalization by randomly skipping the residual branch during training, thus reducing the depth on a per-sample basis.
@@ C.2 Software
For the development of BioAnalyst we used the Python programming language [58] and developed the required modules on PyTorch’s neural network library [59] and PyTorch Lightning [60]. For visualisation we employed the Streamlit package [61]. For data operations we used xarray package [62].
@@ C.3 Training and Scaling
BioAnalyst has been trained for 1000 epochs - 80.000 gradient steps with a batch size of 1, using the AdamW optimiser [63] with cosine-annealing learning rate schedule with periodic warm restarts [64]. We used a starting learning rate of $0.00005$ and weight decay of $0.000005$ with $T_{periodic}=8000$ gradient steps.
@@
To balance the loss during pre-training and subsequent finetuning tasks we assign individual weights to each variable for every variable group. Our weighting scheme is inspired by the works of [18, 65, 66, 67, 68, 69, 70] and are reported on Table 8.
Table 8: Variable-specific weights $(w_{v})$ used in the loss function.
@@
Disclaimer
The results on this version of the manuscript, are produced with the Small model, while we strive to finalise training the Medium model.
C.4 Metrics
C.4.1 Pre-training
@@
$\text{MAE}=\frac{1}{V}\sum_{v=1}^{V}\frac{1}{H\times W}\sum_{i=1}^{H}\sum_{j=1}^{W}||\mathbf{\hat{X}}_{i,j}^{v}-\mathbf{X}_{i,j}^{v}||$ (15)
@@
C.4.2 Task-specific finetuning
@@
[…]
@@
[…]
@@
where $c_{ij}$ is the number of species present in both ground-truth and prediction at cell $(i,j),b_{ij}$ those present only in the observation, and $c_{ij}$ only in the prediction. The mean assemblage similarity estimate is an informative indicator for the performance of the species distribution model (SDM) [71].
Appendix D Further Results
Figure 11: A scorecard highlighting the MAE from the autoregressive rollout (12 steps) for the species variable group

### v2

@@ B.4 Fine-tune datasets
GeoLifeCLEF24: GeoLifeCLEF is benchmark dataset with a training set of close to 5 million plant occurrences in Europe (single-label, presence-only data) as well as a validation set of about 5000 plots and a test set with 20000 plots, with all the present species (multi-label, presence-absence data) (geolifeclef-2024) from 2017 to 2021. A visual depiction of how the data look like can be found on Figure 8.
@@
CHELSA v2.1: CHELSA (Climatologies at High resolution for the Earth’s Land Surface Areas) delivers global grids of mean, minimum and maximum 2 m air temperature (tas, tasmin, tasmax) and precipitation rate (pr) at 30-arc-second resolution ($\sim 1$ km) from 1979 - present (we use 2000-2019 monthly records). Temperature fields are produced through lapse-rate–based statistical downscaling of ERA-5/ERA-Interim reanalyses, while precipitation is downscaled with an orographic algorithm that incorporates wind fields, boundary-layer height and subsequent bias-correction with Global Precipitation Climatology Centre (GPCC) gauge data, yielding markedly improved representation of mountain rain-shadows and thermal gradients compared with coarser products (karger2017climatologies). These high-fidelity temperature and precipitation layers form the foundation for CHELSA’s 19 bioclimatic derivatives and for time-series forcings such as CHELSA-W5E5, and are now a de-facto standard in species-distribution modelling, trait–environment analyses, hydrological and vegetation-dynamics models, and climate-change impact assessments. Because ecological responses to climate are often threshold-driven and spatially heterogeneous, the $\sim 1$ km detail of CHELSA allows modellers to resolve local refugia, elevational turnover and fine-scale moisture stress that coarser climatologies fail to capture, thereby increasing predictive accuracy and reducing uncertainty across taxa and regions (karger2020high).
@@ C.1 Model card
Table 6: Model card
model patch_size num_heads embed_dim depth swin_size model size
Medium 4 12 384 6 medium 440M
Large 8 16 512 10 large 700M
Table 7: Swin-backbone configurations for large and medium model sizes
@@
drop_rate 0.1 0.0
attn_drop_rate 0.1 0.0
@@
Furthermore, to enhance performance and computational efficiency, particularly given the high dimensionality of the input and output data, several techniques are used. Group-Query Attention (ainslie2023gqa) is used within the Perceiver IO attention layers to reduce the memory footprint associated with key-value projections during cross-attention. Additionally, standard regularisation techniques such as Layer Normalisation and Dropout are applied throughout the Perceiver modules. The Swin Transformer backbone makes use of stochastic depth (huang2016deep) to improve generalisation by randomly skipping the residual branch during training, thus reducing the depth on a per-sample basis.
@@ C.2 Software
For the development of BioAnalyst we used the Python programming language (python3.12.0doc) and developed the required modules on PyTorch’s neural network library (paszke2019pytorchimperativestylehighperformance) and PyTorch Lightning (Falcon_PyTorch_Lightning_2019). For visualisation we employed the Streamlit package (streamlit2025). For data operations we used xarray package (hoyer2017xarray).
@@ C.3 Training and Scaling
BioAnalyst has been trained for 1000 epochs - 80.000 gradient steps with a batch size of 1, using the AdamW optimiser (loshchilov2019decoupledweightdecayregularization) with cosine-annealing learning rate schedule with periodic warm restarts (loshchilov2017sgdr). We used a starting learning rate of $0.00005$ and weight decay of $0.000005$ with $T_{periodic}=8000$ gradient steps. We train the complete architecture together, both encoder-decoder and backbone.
@@
To balance the loss during pre-training and subsequent finetuning tasks we assign individual weights to each variable for every variable group. Our weighting scheme is inspired by the works of (bodnar2024aurora; piedallu2013soil; zomer2022version; carbognani2012influence; coelho2023geography; maharjan2022temperature; bates2025soil) and are reported on Table 8.
Table 8: Variable-specific weights $(w_{v})$ used in the loss function
@@
C.4 Finetuning tasks
C.4.1 Roll-out finetuning
BioAnalyst has been pre-trained to predict $x_{t+1}$ by providing as input $x_{t-1},x_{t}]$. During fine-tuning and more specific, roll-out fine-tuning, BioAnalyst was extended to predict 6 and 12 steps ahead in 2 stages. At the first stage, we fine-tune the complete BioAnalyst, adding VeRA adapters on the backbone for 12000 steps with a horizon of 6 time-steps (6 months), using constant learning rate of $0.00003$, weight decay of $0.000003$ and store the weights. In the second stage we load the previously stored weights and continue the roll-out fine-tuning for 12000 time-steps more, with a horizon of 12 steps (1 year), using constant learning rate of $0.00001$ and weight decay of $0.000001$. Fine-tuning with a mix of variable horizons is a technique used on Aurora (bodnar2024aurora), PrithviWxC (schmude2024prithviwxcfoundationmodel) and Aardvak (allen2025end), offering improved long-horizon forecasting capabilities, which we also observe ourselves and are highlighted on Figure 3.
C.4.2 Abiotic linear probing: recovering seasonal climate structure
C.4.3 Biotic fine-tuning: forecasting species distributions
The model takes as input a species distribution matrix at time $t$ and is trained to predict the distribution at time $t+1$. The input and target distributions are normalised using species-specific statistics (mean and standard deviation). The model preserves the full spatial resolution of the data, working with grid-based representations rather than individual occurrence points. Training is performed end-to-end using a combination of loss functions, including the GeoLifeClef defined F1 score (picek2024overview) and root mean square error (RMSE). The custom F1 score is determined in Equation 20. These metrics indicate whether the model learns both local and global distribution patterns, and also gives a comparable score for the GeoLifeCLEF benchmark.
C.4.4 Baselines
We include Aurora-$0.25^{\circ}$ as a strong, climate-only baseline (bodnar2024aurora) by attaching a simple prediction head and fine-tuning it on our task. This comparison is necessarily imperfect: BioAnalyst is pre-trained directly on monthly biodiversity states and multimodal ecological covariates, whereas Aurora is an Earth-system foundation model trained on over a million hours of geophysical data and optimised for short-lead (6-hour) forecasting of atmospheric and oceanic variables, then fine-tuned for specific operational weather, wave and air-quality tasks. As a result, our Aurora baseline should be viewed as an indicative demonstration of how a generic climate foundation model performs when naively adapted to biodiversity prediction, rather than a definitive assessment of Aurora’s capabilities. Our empirical evidence is limited to a small number of downstream biodiversity tasks and one region, and we therefore refrain from drawing general conclusions about Aurora itself.
Joint Species Distribution Modelling
To establish strong yet interpretable baselines for joint spatio–temporal species distribution modelling (jSDM), we consider two complementary architectures. First, we implement a non-spatial latent-factor jSDM that treats each grid cell independently and models the joint community vector with a shared multilayer perceptron (MLP) - LatentMLP. This design follows the generalized linear latent-variable / jSDM tradition (Warton2015; Ovaskainen2017), in which cross-species residual covariance is captured via low-dimensional latent factors rather than separate single-species models. In practice, we encode the presence–absence vector of all species at time $t$ into a low-dimensional community embedding and decode it into species-specific logits for time t+1t{+}1, yielding a multi-output Bernoulli model closely aligned with recent scalable jSDM implementations such as sjSDM (PichlerHartig2021). This baseline provides a conceptually clean joint model that captures species co-occurrence structure but ignores explicit spatial autocorrelation.
Second, we introduce a spatio–temporal ConvLSTM jSDM that operates directly on the $C\times H\times W$ species distribution grids over consecutive years. We first project the species axis into a smaller set of learned community channels using $1{\times}1$ convolutions, then apply a convolutional LSTM over the temporal dimension to propagate information across both space and time before decoding back to species-level probabilities with a sigmoid output. ConvLSTMs were originally proposed for precipitation nowcasting as powerful spatio–temporal sequence models (Shi2015) and are now widely used in environmental forecasting. Our baseline adapts this idea to multi-species distributions, in line with recent work showing that joint deep neural networks improve community-level predictions compared to species-wise models (Brun2024) and that deep-SDM frameworks such as MALPOLON scale to thousands of species and high-resolution remote-sensing inputs (Larcher2024). Finally, by training and evaluating these baselines on GeoLifeCLEF2024 presence–absence grids, we align our setup with current large-scale benchmarks for species composition and presence prediction that explicitly target continental-scale, high-resolution community modelling (picek2024overview; Picek2025GeoLifeCLEF). Together, the LatentMLP and ConvLSTM baselines span a representative spectrum from classical joint SDMs to modern spatio–temporal deep jSDMs, providing a strong baselines to compare BioAnalyst.
Climate linear probing.
To quantify how much climate-relevant information is encoded in our learned representations, we adopt a climate linear probing setup: the encoder is frozen and we fit a single linear layer to predict gridded climate variables or indices from the embeddings, following the idea of linear classifier probes as a measure of linear separability in deep features (alain2016understanding). To contextualise these linear-probe scores against widely used non-parametric methods in climate science, we additionally train two baselines on the same features: a Random Forest (RF) and a Support Vector Machine (SVM). RFs have recently been shown to be highly effective for statistical downscaling and bias adjustment of climate simulations over Europe, for example through a posteriori random forests that stochastically downscale daily precipitation at dozens of European stations while accurately capturing local conditional distributions (legasa2022posteriori). Complementarily, SVM-based approaches have been used at the global scale to post-process subseasonal-to-seasonal precipitation forecasts from the European Centre for Medium-Range Weather Forecasts, significantly improving skill over the raw dynamical model output (yin2023improving). Reporting linear-probe performance alongside RF and SVM therefore allows us to disentangle how much climate signal is linearly accessible in the representation space of BioAnalyst versus what can be extracted by strong off-the-shelf non-linear baselines.
C.5 Metrics
C.5.1 Pre-training
@@
[…]
@@
C.5.2 Task-specific finetuning
@@
where $\mathrm{TP}_{i}$ $=\text{number of correctly predicted species (true positives)},$
$\mathrm{FP}_{i}$ $=\text{species predicted but \emph{not} observed (false positives)},$
$\mathrm{FN}_{i}$ $=\text{species present but \emph{not} predicted (false negatives)},$
@@
$R^{2}\;=\;1-\frac{\displaystyle\sum_{i=1}^{n}\bigl(y_{i}-\hat{y}_{i}\bigr)^{2}}{\displaystyle\sum_{i=1}^{n}\bigl(y_{i}-\bar{y}\bigr)^{2}},\qquad\bar{y}\;=\;\frac{1}{n}\sum_{i=1}^{n}y_{i},$ (21)
@@
where $c_{ij}$ is the number of species present in both ground-truth and prediction at cell $(i,j),b_{ij}$ those present only in the observation GBIF data, and $d_{ij}$ species only predicted by the model. The mean assemblage similarity estimate is an informative indicator for the performance of the species distribution model (SDM) (pinto2021predicting).
Our community comparison uses an incidence-based Sørensen coefficient per grid cell, the similarity pattern captures a combination of (i) actual model skill and (ii) where the GBIF data are dense and reliable enough to evaluate species assemblages. In well-sampled regions, more species are recorded, detection probabilities are higher, and the model is trained on richer data; not surprisingly, these are also the areas where Sørensen similarity tends to be highest. In poorly sampled regions, many species are undetected, so unobserved but present species are treated as absences in the metric, which depresses Sørensen scores regardless of how good the underlying model might be. Assemblage-level SDM evaluations using Sørensen and related indices are known to be sensitive to such incompleteness. In addition, GBIF data are known to be geographically and taxonomically biased  [^1] https://www.gbif.org/data-use/6hL40kh9ikDXftobIM85KP/sampling-biases-shape-our-view-of-the-natural-world, with much denser sampling in Western and Central Europe, near cities and easily accessible areas, and substantially sparser coverage in Eastern and South-eastern Europe and remote regions. This pattern has been documented repeatedly for GBIF and similar occurrence databases (beck2014spatial).
Appendix D BioAnalyst ablations and interpretability
In order to understand which input modalities contribute most to biodiversity predictions, we decided to observe the cross-attention weights from the encoder’s first layer, which directly measures how the model allocates computational focus across input features before hierarchical processing in the Swin backbone. For a representative test window (December 2019), consider Figure 9, which shows that climate variables receive the highest attention, followed by species and surface observations, while most other modality groups receive lower attention. This ranking seems consistent across the 13 prediction windows in the test set Figure 10), with climate, species and surface maintaining high attention. Spatial analysis of the same representative window, particularly in the case of a high-interest variable (i.e., species), reveals that attention is not uniformly distributed: the encoder concentrates on Scandinavia and Central Europe regions Figure 12), which are hotspots where the strongest species occurrence signals would be encountered in the used data set. We focused on encoder cross-attention rather than decoder attention because it provides a direct, interpretable measure of which input modalities the model prioritises during initial feature extraction, before representations are hierarchically mixed by the Swin backbone, making modality contributions easier to quantify and attribute.
Additionally, Table LABEL:tab:modality_ablation demonstrates the MAE performance across different variable groups, over the test set. The columns represent specific modality ablation studies: C1 (Species $\times$ Climate), C2 (Species $\times$ Edaphic), C3 (Species $\times$ RLI), and C4 (Species $\times$ Surface), compared against a more complex model C5 (Species $\times$ Edaphic $\times$ RLI $\times$ Climate $\times$ Surface) and the BFM Full model combining all listed previously variables. Cells marked with ”XXX” indicate variables that are out of scope for that specific modality combination. The results show that while the more comprehensive models (C5 and BFM Full) provide consistent predictions across physical variables (Surface, Edaphic, and Climate groups), single-modality combinations can offer slightly better performance in specific tasks. For example, for species abundance predictions (bottom section), the Climate-only model (C1) frequently outperforms the fuller models (e.g., for species ID 9809229, C1 achieves an MAE of 0.627 compared to 0.740 for C5). This could suggest that either climatic variables are the dominant helpers in predicting species distributions in this dataset, and adding additional modalities (as in C5) may occasionally introduce noise rather than distinct predictive signals for certain species.
Figure 9: Mean attention weight per modality group contribution for the Encoders cross-attention layer.
Figure 10: Mean attention weight per modality group contributions for 12 test windows (months).
Figure 11: Cross-modality attention correlation matrix for all the modality groups with their respective variables.
Figure 12: Cross-modality spatial attention map for the species variable group.
Table 9: Performance comparison between different modality combinations
Group $(V)$ Var $(v)$ C1 C2 C3 C4 C5 BFM Full
Surface t2m XXX XXX XXX 2.753380 0.003875 3.007635
msl XXX XXX XXX 433.657866 421.848398 400.324847
slt XXX XXX XXX 0.005178 0.008991 0.013444
z XXX XXX XXX 16.460119 23.041222 33.079982
u10 XXX XXX XXX 1.344401 1.260686 1.218697
v10 XXX XXX XXX 1.156675 1.093683 1.105192
lsm XXX XXX XXX 0.002361 0.003875 0.004778
Edaphic swvl1 XXX 0.143530 XXX XXX 0.015753 0.016012
swvl2 XXX 0.139144 XXX XXX 0.016992 0.016745
stl1 XXX 6.284586 XXX XXX 2.444646 2.544988
stl2 XXX 6.185118 XXX XXX 2.403322 2.490814
Atmospheric (p-levels) z_pl XXX XXX XXX XXX XXX 77295.114183
t_pl XXX XXX XXX XXX XXX 39.698364
u_pl XXX XXX XXX XXX XXX 6.898406
v_pl XXX XXX XXX XXX XXX 1.645421
q_pl 0.78 XXX XXX XXX XXX 0.005081
Climate smlt 0.000228 XXX XXX XXX 0.000116 0.000124
tp 0.001278 XXX XXX XXX 0.000890 0.000857
csfr 0.000001 XXX XXX XXX 0.000001 0.000001
avg_sdswrf 80.703410 XXX XXX XXX 35.588897 34.508076
avg_snswrf 71.948338 XXX XXX XXX 31.350053 30.553276
avg_snlwrf 15.075831 XXX XXX XXX 7.237449 7.593424
avg_tprate 0.000015 XXX XXX XXX 0.000010 0.000010
avg_sdswrfcs 98.768588 XXX XXX XXX 44.961979 44.932251
sd 0.060704 XXX XXX XXX 0.031556 0.018263
t2m_clim 6.429481 XXX XXX XXX 2.730068 3.021820
d2m 5.719590 XXX XXX XXX 2.675545 2.732889
Vegetation NDVI XXX XXX XXX XXX XXX 0.033334
Land cover Land XXX XXX XXX XXX XXX 982.858652
Agriculture Agriculture XXX XXX XXX XXX XXX 0.369489
Arable XXX XXX XXX XXX XXX 0.196134
Cropland XXX XXX XXX XXX XXX 0.035131
Forest Forest XXX XXX XXX XXX XXX 0.201663
Redlist RLI XXX XXX 0.002770 XXX 0.003895 0.004125
Miscellaneous avg_slhtf XXX XXX XXX XXX XXX 5.487877
avg_pevr XXX XXX XXX XXX XXX 0.000004
Species 1340361 0.000042 0.000061 0.000042 0.000046 0.000043 0.000064
1340503 0.063857 0.064053 0.063795 0.063897 0.063929 0.064850
1536449 0.034173 0.037600 0.034158 0.034184 0.034193 0.034896
1898286 0.322074 0.383199 0.335297 0.356860 0.358814 0.359776
1920506 0.276737 0.267358 0.244421 0.264019 0.265405 0.272244
2430567 0.001140 0.009217 0.001117 0.001168 0.001171 0.001872
2431885 0.011713 0.012341 0.011626 0.011719 0.011766 0.012577
2433433 0.000929 0.000936 0.000929 0.000934 0.000934 0.001001
2434779 0.000014 0.000021 0.000014 0.000014 0.000014 0.000018
2435240 0.004547 0.007200 0.004532 0.004726 0.004585 0.005042
2435261 0.000050 0.000066 0.000049 0.000052 0.000052 0.000074
2437394 0.000181 0.000183 0.000180 0.000183 0.000184 0.000220
2441454 0.004198 0.004282 0.004192 0.004248 0.004272 0.005205
2473958 0.049975 0.046870 0.049680 0.049740 0.049872 0.051274
2491534 0.246025 0.252939 0.267842 0.282393 0.284369 0.294113
2891770 0.043579 0.043935 0.043551 0.043735 0.043834 0.046475
3034825 0.021265 0.021283 0.021218 0.021394 0.021357 0.023135
4408498 0.015352 0.015404 0.015298 0.015445 0.015424 0.016373
5218786 0.007418 0.007418 0.007412 0.007432 0.007430 0.007713
5219073 0.000768 0.000766 0.000763 0.000770 0.000769 0.000831
5219173 0.006217 0.006225 0.006212 0.006230 0.006233 0.006481
5219219 0.000254 0.000255 0.000254 0.000261 0.000256 0.000296
5844449 0.006492 0.006508 0.006476 0.006517 0.006520 0.006944
8002952 0.015073 0.015116 0.015069 0.015131 0.015135 0.015838
8077224 0.318140 0.329542 0.340779 0.359972 0.357112 0.361090
8894817 0.000202 0.000228 0.000199 0.000217 0.000208 0.000336
8909809 0.005668 0.005821 0.005664 0.005761 0.005711 0.006676
9809229 0.627497 0.646721 0.719715 0.735913 0.739663 0.745430